OpenStudy (anonymous):

cosec2(90−θ)+cos2(90−θ)−cot245 please help

OpenStudy (anonymous):

\[cosec^{2}(90−θ)+\cos^{2}(90-\theta)−\cot^{2}45\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

looks up the "cofuntion identities" to get rid of the \(90-\theta\) part

OpenStudy (anonymous):

\[cosec^{2}(90 degree−θ)+\cos^{2}(90degree −θ)−\cot^{2}45 degree\]

OpenStudy (anonymous):

please find the value

OpenStudy (anonymous):

sec^{2}90 degree+sin^{2}90 degree-cot^{2} 45 degree

OpenStudy (anonymous):

0+1-1